OpenStudy (anonymous):

Shanna writes the formula f(x + 1) = 2.5f(x) when f(1) = 2 to represent this sequence: 2, 5, 12.5, 31.25, … Which error did Shanna make?

OpenStudy (anonymous):

A) She used the incorrect common ratio. B) She used the incorrect initial value. C) She should have multiplied by f(x) rather then have it as an exponent. D) She treated the sequence as geometric instead of arithmetic.

OpenStudy (owlcoffee):

Well, you'll have to determine the common ratio the same way, divide the second by the first, and if it does not give 2,5 as a result, she used the wrong common ratio.

OpenStudy (anonymous):

what do you think @shaylaLuv ?

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

I think shanna did nothing wrong

OpenStudy (anonymous):

\(f(x + 1) = 2.5f(x)\), \(f(1) = 2\), let \(x=1\)\[f(2)=2.5f(1)=2.5 \times 2=5\]and let \(x=2\)\[f(3)=2.5f(2)=2.5 \times5=12.5\]let \(x=3\)\[f(4)=2.5f(3)=2.5 \times 12.5=31.25\]